ROTATIONAL SPECTRA OF HALOGENATED ETHERS USED AS VOLATILE ANAESTHETICS

ALICIA VEGA-TORIBIO and \underline{ALBERTO LESARRI}, Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47011 Valladolid, Spain; RICHARD D. SUENRAM, Department of Chemistry, University of Virginia, Charlottesville, VA 22904, USA; JENS-UWE GRABOW, Institut für Physikalische Chemie und Elektrochemie, Leibniz Universität Hannover, Callinstraße 3 - 3A, D-30167 Hannover, Germany.

Following previous microwave investigations by Suenram et al.^{a} we will report on the rotational spectrum of several halogenated ethers used as volatile anaesthetics, including sevoflurane ((CF_{3})_{2}CH-O-CH_{2}F), isoflurane (CF_{3}CHCl-O-CHF_{2}), enflurane (CHFClCF_{2}-O-CHF_{2}) and methoxyflurane (CHCl_{2}CF_{2}-O-CH_{3}). This study has been conducted in the 6-18 GHz centimetre-wave region using Balle-Flygare-type FT-microwave spectroscopy. The results will include the analysis of the rotational spectra of minor species in natural abundance (^{13}C and ^{18}O in some cases), structural calculations and auxiliary ab initio modelling. The conformational and structural conclusions will be compared with previous gas-phase electron diffraction and solid-state X-ray diffraction analysis.

aR. D. Suenram, D. J. Brugh, F. J. Lovas and C. Chu, 51^{st} OSU Int. Symp. On Mol. Spectrosc., Columbus, OH, 1999, RB07